OpenStudy (vortish):

solve the system by the substitution method x+y=3 y=x^2-6x+9

OpenStudy (anonymous):

okay just plug in the second equation into the first equaiton

OpenStudy (vortish):

so x+x^2-6x+9=3 right then you subtract 3 from both sides right so you get x^2+7x-12=0 ?

OpenStudy (anonymous):

i would do something different.... substitution is all about plugging in so the first equation looks the easiest to start with.. so let's just isolate the x :) so we have x= -y+3 now, you substitute it into that second equation to find y so y=(-y+3)^2-6(-y+3)+9 and now solve piece by piece (-y+3)(-y+3) use FOIL to solve so y^2-3y-3y+9 = y^2-6y+9 then you have -6 times -y = 6y and then -18 so which gives you y^2-6y+9+6y-18+9. combine like terms! what do you get?

OpenStudy (vortish):

this is non linear system substittuion

OpenStudy (anonymous):

ohh okay so then never mind hahah do what the first person said :) haha what you got looks right :)

OpenStudy (anonymous):

and then continue from there! :D